# The building blocks of software work explain coding careers and language popularity

### v1

@@ D.3 Regression of entry on task value
Detailed outcomes of the regression model described in eq. (9) of the main text. Observations are user-task combinations for which the user did not answer any questions related to the task in the previous two years. The dependent variable is a binary value of whether or not the user enters the task (i.e., answers questions related to this task) in the current period. task value: logarithm of the task’s imputed value, task density: the user’s prior experience in other tasks, weighted by the relatedness of these tasks to the current task. For ease of interpretation, we subtract the variable’s mean and then divide by its standard deviation. Models in columns (2), (4), (6) and (8) control for user fixed effect. Columns (1)-(4) report analyses performed using all answer posts, columns (5)-(8) use only Python related answer posts. Clustered (user) standard-errors in parentheses. p-values: ***: 0.01, **: 0.05, *: 0.1.
@@ Appendix E Rescaling by Github
Figure 4 shows Python’s progressive rise as the dominant language for an increasing number of programming tasks. However, even though SO may provide a useful dataset to define such programming tasks, it is not guaranteed to be representative of programmers across the world. To assess this problem, we compare counts of programmers on SO to comparable ones on GitHub. GitHub is the world’s largest software development collaboration platform and therefore offers a good representation of the global software development sector. In particular, we collect information on the (self-reported) countries of residence on GitHub and (observed) language use between 2020 and 2022 [GitHub, 2025]. Figure E.1a compares SO to GitHub user counts at the level of countries, languages and country-language combinations. This shows that user numbers are highly correlated across the two platforms, with correlations ranging from 0.75 to 0.93.
Figure E.1: a. The GitHub user counts against the SO user counts in the perspectives of country, programming languages and country-language in the year 2022. The values are in log scale and the $R^{2}$ values of the regression fit are provided inside each panel. b. Number of tasks in which a programming language is the top language in terms of SO users reweighted by user counts from GitHub, from 2020 to 2022. The graph shows the largest 6 languages measured by cumulative SO posts between August 2008 and June 2023.
Nevertheless, some languages are better represented on SO than others. This may affect our analysis of how dominant a language is for a given task. In Fig. E.1b we attempt to correct for this, by reweighting the user counts by language using their observed shares on GitHub. This reweighting ensures that the distribution of users across language on SO mimics the one on GitHub. Because we only have GitHub information for the years 2020, 2021 and 2022, the analysis is limited to those years. Also with reweighted user counts, Python leads in most tasks, albeit with round 50, in somewhat fewer tasks than what we report in the main text. The ordering of the remaining languages is mostly unchanged as well, except for C#, which was somewhat more prominent before rescaling user counts. Moreover, also the most prominent temporal patterns over the 3-year time period are mostly left unchanged. Figure E.1b retains the rise of Python and the drops of Java and PHP, as well as the relative stability in these years of JavaScript, validating the results shown in the main text.
@@ Appendix F Swift vs Objective-C
Figure F.1: Number of tasks in which Objective-C has more users on SO than Swift and vice versa. The dashed vertical line marks the year 2014, the year that Swift was released.
@@
Figure F.1 describes this shift in terms of the tasks where Swift was more popular than Objective-C and vice versa. Before 2015, most tasks were handled preferentially in Objective-C. However, since the release of Swift in 2014 this rapidly changes. Swift surpasses Objective-C in terms of the number of different tasks it handles around 2015 and all but eclipses Objective-C by 2016.

### v2

@@ D.3 Regression of skill acquisition on skill value
Detailed outcomes of the regression model described in eq. (9) of the main text. Observations are user-skill combinations for which the user did not answer any questions related to the skill in the previous two years. The dependent variable is a binary variable that encodes whether or not the user acquires the skill (i.e., answers questions related to this skill) in the current period. skill value: logarithm of the skill’s imputed value, skill density: the user’s prior expertise in other skills, weighted by the relatedness of these skills to the current skill. For ease of interpretation, we subtract the variable’s mean and then divide by its standard deviation. Models in columns (2), (4), (6) and (8) control for user fixed effect. Columns (1)-(4) report analyses performed using all answer posts, columns (5)-(8) use only Python related answer posts. Clustered (user) standard-errors in parentheses. p-values: ***: 0.01, **: 0.05, *: 0.1.
@@ Appendix E Rescaling by Github
Fig. 3 shows Python’s progressive rise as the dominant language used in an increasing number of programming skills. However, even though SO may provide a useful dataset to define such programming skills, it is not guaranteed to be representative of programmers across the world. To assess this problem, we compare counts of programmers on SO to comparable ones on GitHub. GitHub is the world’s largest software development collaboration platform and therefore offers a good representation of the global software development sector. In particular, we collect information on the (self-reported) countries of residence on GitHub and (observed) language use between 2020 and 2022 [GitHub, 2025]. Fig. E.1a compares SO to GitHub user counts at the level of countries, languages and country-language combinations. This shows that user numbers are highly correlated across the two platforms, with correlations ranging from 0.75 to 0.93.
Figure E.1: a. The GitHub user counts against the SO user counts in the perspectives of country, programming languages and country-language in the year 2022. The values are in log scale and the $R^{2}$ values of the regression fit are provided inside each panel. b. Number of s in which a programming language is the top language in terms of SO users reweighted by user counts from GitHub, from 2020 to 2022. The graph shows the largest 6 languages measured by cumulative SO posts between August 2008 and June 2023.
Nevertheless, some languages are better represented on SO than others. This may affect our analysis of how dominant a language is in a given skill. In Fig. E.1b we attempt to correct for this, by reweighting the user counts by language using their observed shares on GitHub. This reweighting ensures that the distribution of users across language on SO mimics the one on GitHub. Because we only have GitHub information for the years 2020, 2021 and 2022, the analysis is limited to those years. Also with reweighted user counts, Python leads in most skills, albeit with around 50, in somewhat fewer skills than what we report in the main text. The ordering of the remaining languages also remains mostly unchanged, except for C#, which was somewhat more prominent before rescaling user counts. Moreover, also the most prominent temporal patterns over the 3-year time period are mostly left unchanged. Fig. E.1b retains the rise of Python and the drops of Java and PHP, as well as the relative stability in these years of JavaScript, validating the results shown in the main text.
@@ Appendix F Swift vs Objective-C
Figure F.1: Number of skills in which Objective-C has more users on SO than Swift and vice versa. The dashed vertical line marks the year 2014, the year that Swift was released.
@@
Fig. F.1 describes this shift in terms of the skills where Swift was more popular than Objective-C and vice versa. Before 2015, most skills were deployed preferentially using Objective-C. However, since the release of Swift in 2014 this rapidly changes. Swift surpasses Objective-C in terms of the number of different skills in which it dominates around 2015 and all but eclipses Objective-C by 2016.